OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through point (1, -2) and has a slope of 1/3. 3x - y - 7 = 0 x - 3y + 7 = 0 x - 3y - 7 = 0

OpenStudy (nory):

First write it in point-slope form, and then rearrange terms.

OpenStudy (nory):

y+2 = 1/3(x-1) is point-slope form. Now rearrange the terms.

OpenStudy (anonymous):

Point Slope form being the following, place your points into point slope and move everything over into general form. Y=mx+b \[Y-Y _{1}=m(X-X _{1})\]